OpenStudy (anonymous):

Solve the system by substitution. -4.5x-2y= -12.5 3.25x-y= -0.75

OpenStudy (alekos):

multiply the bottom equation by -2 and then add the two equations

OpenStudy (anonymous):

So -2(3.25x-y= -0.75)

OpenStudy (anonymous):

?

OpenStudy (alekos):

now multiply that out

OpenStudy (anonymous):

Okay give me a minute.

OpenStudy (anonymous):

I can't do it on my calculator because i cant put in the = sign.

OpenStudy (anonymous):

-6.5-2y=1.5

OpenStudy (alekos):

forget the calculator. use a pen and paper

OpenStudy (anonymous):

Did i get it right?

OpenStudy (alekos):

its +2y

OpenStudy (anonymous):

Oh your right. Two negatives make a positive.

OpenStudy (anonymous):

So now what do i do?

OpenStudy (alekos):

Now add the two equations term by term and put your answer under the dotted line -4.5x - 2y = -12.5 -6.5x + 2y = 1.5 ------------------

OpenStudy (anonymous):

So i would add -4.5x and -6.5x.

OpenStudy (alekos):

yes and then the rest

OpenStudy (anonymous):

-11

OpenStudy (anonymous):

-2y+2y=0

OpenStudy (alekos):

you need to write that as one new equation underneath the dotted line term by term

OpenStudy (anonymous):

-12.5+1.5=-11

OpenStudy (anonymous):

-4.5x - 2y = -12.5 -6.5x + 2y = 1.5 ------------------ -11 + 0 = -11

OpenStudy (alekos):

first term is -11x

OpenStudy (anonymous):

if you need more help ask @xMissAlyCatx

OpenStudy (alekos):

so we get a new equation of -11x=-11

OpenStudy (anonymous):

Okay

OpenStudy (alekos):

go away gravyboat

OpenStudy (alekos):

now we divide both sides by -11 and we have a value for x

OpenStudy (alekos):

I'll let you do that

OpenStudy (anonymous):

1x?

OpenStudy (alekos):

No, try again

OpenStudy (alekos):

Gotta go out for dinner so I'll get back to you in a couple of hors

OpenStudy (anonymous):

Oh okay thanks.